OpenStudy (anonymous):

what is the solution of (3x+4)=14?

OpenStudy (mathstudent55):

First, drop the parentheses on the left side. They are not needed.

OpenStudy (anonymous):

Okay so 3x+4=14? do I subtract the 4? first?

OpenStudy (mathstudent55):

You want x alone on the left side. If you wanted to evaluate the expression 3x + 4 for a value of x, you'd first multiply that value of x by 3. Then you'd add 4 to the result. To get x alone, you need to do the opposite operations in the opposite order. The opposite of add 4 is subtract 4, so subtract 4 from both sides.

OpenStudy (mathstudent55):

Then you divide both sides by 3 since division by 3 is the opposite operation of multiplying by 3.

OpenStudy (anonymous):

So X=3?

OpenStudy (anonymous):

x=10.7 because when u devide everything u add it all together to see if it plugs back in to equal 14

OpenStudy (anonymous):

and then x= 3.3

OpenStudy (mathstudent55):

3x + 4 = 14 Subtract 4 from both sides: 3x + 4 = 14 -4 -4 ---------- 3x = 10 Now divide both sides by 3: 3x/3 = 10/3 Solution is x = 10/3

OpenStudy (anonymous):

or 3.3 its the same thing

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

welcome can i get a medal

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

how do i give you a medal? i'm new to this

OpenStudy (mathstudent55):

Click on Best Response in one of my answers.